OpenStudy (anonymous):

prove that 2^n>=n^2 for n>=4

OpenStudy (zarkon):

use mathematical induction

OpenStudy (anonymous):

i am at the point that 2^(k+1)>=(k+1)^2

OpenStudy (anonymous):

than i multiplied by 2 the two sides

OpenStudy (anonymous):

of this 2^k>=k^2

OpenStudy (anonymous):

and i got 2^(k+1)>=2k^2

OpenStudy (anonymous):

Start off with n = 4 2^4 = 4^2 GOOD For n = 5 2^5 > 5^2 GOOD Now assume that it holds for n, i.e. 2^n ≥ n^2 Let's try n+1 2^(n+1) = 2*2^n ≥ 2*n^2 = n^2 + n^2 Now since n ≥ 4, n^2 ≥ 4n = 2n + 2n and 2n > 1 So n^2 > 2n + 1 Thus 2n^2 > n^2 + 2n + 1 = (n+1)^2 Thus 2^(n+1) = 2*2^n > (n+1)^2

OpenStudy (ranga):

2^n >= n^2 for n >= 4 When n = 4, 2^n = 2^4 = 16; n^2 = 4^2 = 16. So 2^n >= n^2 is true. When n = 5, 2^n = 2^5 = 32; n^2 = 5^2 = 25. So 2^n >= n^2 is true. When n = k, 2^n = 2^k; n^2 = k^2. Assume 2^k >= k^2. With the assumption 2^k >= k^2 if we can prove that 2^(k+1) >= (k+1)^2 then we have proved the original question by induction. If 2^k >= k^2, then: 2 * 2^k >= 2 * k^2 2^(k+1) >= 2k^2 ---- (1) Now we need to prove 2k^2 >= (k+1)^2: 2k^2 >= k^2 + 2k + 1 k^2 - 2k >= 1 k^2 - 2k + 1 >= 2 (k-1)^2 >= 2 This is indeed true for all values of k > 2. The problem is only interested in k >= 4. Therefore, 2k^2 >= (k+1)^2 for k >= 4. Put this in (1): 2^(k+1) >= (k+1)^2 for k >= 4 We started with the assumption that 2^k >= k^2 and proved that 2^(k+1) >= (k+1)^2. Therefore, by induction we have proved that 2^n >= n^2 for n >= 4.